OpenStudy (braves12):

Solve the following equation by factoring. 6x 2 - 5x = 50 {-5/3, 5} {-5/2, 10/3} {-10/3, 5/2}

OpenStudy (braves12):

@GIitter

OpenStudy (giitter):

@Unavailabilityy

OpenStudy (jango_in_dtown):

\[6x^2-5x-50=0\]

OpenStudy (braves12):

yes this is the probblem

OpenStudy (jango_in_dtown):

6*50= 20 *15

OpenStudy (giitter):

I think i can get the answer hang on

OpenStudy (jango_in_dtown):

and 20-15=5

OpenStudy (jango_in_dtown):

now its easy

OpenStudy (jango_in_dtown):

\[6x^2-20x+15x-50=0\]

OpenStudy (jango_in_dtown):

now factor yourself. the work is done

OpenStudy (giitter):

I got it

OpenStudy (jango_in_dtown):

2x(3x-10)+5(3x-10)=0

OpenStudy (jango_in_dtown):

@braves12 are you doing the problem?

OpenStudy (braves12):

yes

OpenStudy (braves12):

\[x=\frac{ 10 }{ 3 },-\frac{ 5 }{ 2}\]

OpenStudy (jango_in_dtown):

correct

OpenStudy (braves12):

thanks